Radio waves travel at a speed of 1.7 times 10^8 m/
Answer #1

here,

velocity of wave, v = 1.7 * 10^8 m/s
time, t = 32.9*10^-6 s

From first eqn of motion,
acceleration, a = v/t = (1.7 * 10^8)/(32.9*10^-6) = 5.167*10^12 m/s^2

From Third equation of motion, distance ( or depth of ice)
d = v^2/2a
d = (1.7 * 10^8)^2/(2*5.167*10^12)
d = 2796.594 m or 2.796 km

the depth of ice is 2796.594 m